rofessor: #12: SYNTHESIS OF SOAP Date: Section: DA
Can someone answer the quantitative data section with the calculation above. I can't figure it out.
Answer #1

balance the reaction

let reactant R and NaOH and product HOCH2CHOCH2OH (R2) and other (R3)

R + 3 NaOH -----> R2 + 3R3

1 mole of reactant R react with 3 mole of NaOH and product will be

1 mole of HOCH2CHOCH2OH (R2) and 3 mole of (R3) product

moles of reactant (R) is 0.044 moles and 0.0445 moles NaOH.

so 0.0445 moles of NaOH will only react with (0.0445 / 3) moles of reactant R = 0.0148 moles R

and product will be R2 = 0.0148 moles and R3 = 0.0445 moles

theoretical mass of R2 and R3

R2 = (0.0148 moles x 92 g/mol) = 1.3646 g

R3 = (0.0445 moles x 278.2 g/mol) = 12.3799 g

actual mass of R2 = 4.0855 g

and R3 = 12.385 g

percentage yeild R2 = (4.0855 g/1.3646 g) x100 = 299.4 %

percentage yeild R3 = (12.385g/12.3799 g) x100 = 100.04 %